A bullet is fired straight upward with the velocity of 98 m/s from the

top of a building 100 m high. Find: a.) its maximum height above the ground b.) how long is

it in the air? c.) the velocity when it reaches the ground.


1
Expert's answer
2021-05-10T17:23:23-0400

(a) Let's first find the time that the bullet takes to reach its maximum height above the ground:


"v=v_0+gt,""0=v_0+gt,""t=\dfrac{-v_0}{g}=\dfrac{-98\ \dfrac{m}{s}}{-9.8\ \dfrac{m}{s^2}}=10\ s."

Finally, we can find its maximum height above the ground from the kinematic equation:


"y_{max}=y_0+v_0t+\dfrac{1}{2}gt^2,""y_{max}=100\ m+98\ \dfrac{m}{s}\cdot10\ s+\dfrac{1}{2}\cdot(-9.8\ \dfrac{m}{s^2})\cdot(10\ s)^2,""y_{max}=590\ m."

(b) We can find the time of the bullet in the air from the kinematic equation:


"y=y_0+v_0t+\dfrac{1}{2}gt^2,""0=100+98t-4.9t^2,""4.9t^2-98t-100=0."

This quadratic equation has two roots: "t_1=20.97\ s", "t_2=-0.97 s". Since time can't be negative, the correct answer is "t=20.97\ s".

(c) We can find the velocity when bullet reaches the ground from the kinematic equation:


"v=v_0+gt,""v=98\ \dfrac{m}{s}+(-9.8\ \dfrac{m}{s^2})\cdot20.97\ s=-107.5\ \dfrac{m}{s}."

The sign minus means that the velocity of the bullet directed downward.